If two coins, A and B are tossed simultaneously, find the probability of one-coin landing heads up and the other landing tails up in the following order.

a) Probability of coin A landing heads up

b) Probability of coin B landing tails up

c) Probability of coin A landing heads up and coin B tails up

d) Probability of coin A landing tails up

e) Probability of coin A landing tails up and coin B landing heads up



1
Expert's answer
2021-11-15T14:32:29-0500

a)


"P(A\ H)=\dfrac{1}{2}"

b)


"P(B\ T)=\dfrac{1}{2}"

c)


"P(A\ H \& B \ T)=\dfrac{1}{2}(\dfrac{1}{2})=\dfrac{1}{4}"

d)


"P(A\ T)=\dfrac{1}{2}"

e)


"P(A\ T\& B \ H)=\dfrac{1}{2}(\dfrac{1}{2})=\dfrac{1}{4}"